• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name silwal

    Log entry time 06:42:58 on April 06, 2010

    Entry number 316631

    keyword="old" Pb#3 runs @ 50uA

    List of runs taken on Pb#3 target @ 50uA. These were taken on 29Mar10. runs # 2373-2386.

    I updated the pedestals on the .db files for these runs, set the lowbeam cut on curmon bmc3 to be 30000, and reprompted them. I used the pedestal values from bcm/bpm calibration run # 2524.

    We only have bcms 1-4 for these runs. I think we changed a bunch of cable maps into the adcs between these runs and 2524. I checked for map changes briefly, and found that the maps for bcm1s were the same. But these maps should be checked again, if the data does not make sense...



    A copy of this log entry has been emailed to: paschke@jlab.org